At 25 °C, how many dissociated OH− ions are there in 1291 mL of an aqueous solution whose pH is 1.35?
Answer #1

use:

pH = -log [H+]

1.35 = -log [H+]

[H+] = 4.467*10^-2 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(4.467*10^-2)

[OH-] = 2.239*10^-13 M

We have:

volume , V = 1.291*10^3 mL

= 1.291 L

use:

number of mol,

n = Molarity * Volume

= 2.239*10^-13*1.291

= 2.891*10^-13 mol

use:

number of OH- ions = number of moles * Avogadro’s number

= 2.891*10^-13 * 6.022*10^23 OH- ions

= 1.741*10^11 OH- ions

Answer: 1.741*10^11 OH- ions